OpenStudy (anonymous):

Find value of log tan1 + log tan2 + log tan3 + . . .+ logtan 89

OpenStudy (anonymous):

its value is 0 correct? why?

OpenStudy (anonymous):

Hint : \[\log_{} 1 = 0\]

OpenStudy (shubhamsrg):

another hint : tan1 = cot 89 and so on..

OpenStudy (anonymous):

yeah i like that hint better. i guess we were supposed to know we are dealing with degrees.

OpenStudy (anonymous):

*

OpenStudy (anonymous):

\[\log a + \log b = \log (a*b)\]